8. American Airlines flight time from LA to Las Vegas is uniformly distributed between 60 and 70 minutes.

What is the Mean and Variance of the flight times?

Estimate the Probability that flight time is less than 64 minutes.
Answer #1

Here, the given values of lower limit, a = 60 and upper limit, b = 70
For Uniform distribution,
Mean = (a + b)/2
Mean = (60 + 70)/2 = 65

Standard Deviation = sqrt((b - a)^2/12)
Standard Deviation = sqrt((70 - 60)^2/12 = 2.8868

variance = std.deviatio^2
= 2.8868^2
= 8.33


Here, the given values of lower limit, a = 60 and upper limit, b = 70

For Uniform Distribution,
P(X <= x) = (x - a)/(b - a)
P(X <= 64) = (64 - 60)/(70 - 60)
P(X <= 64) = 0.4